The length of a rectangle is 4cm longer than its width. If the perimeter of the rectangle is 40cm,find its area.
Answer by macston(5194)

.
W=width; L=length=W+4cm
.
P=2(L+W)
40cm=2(W+4cm+W)
20cm=2W+4cm
16cm=2W
8cm=W
Width is 8 centimeters.
L=W+4cm=8cm+4cm=12cm
Length is 12 centimeters.
.
A+LW
A=(12cm)(8cm)
A=96cm^2
ANSWER: The area is 96 square centimeters.